Circular No. 6668

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

η CARINAE

K. Ishibashi and K. Davidson, University of Minnesota; M. F. Corcoran, Goddard Space Flight Center (GSFC) and Universities Space Research Association; and J. H. Swank, R. Petre, and K. Jahoda, GSFC, report: “The Proportional Counter Array on board the Rossi X-ray Timing Explorer (RXTE) has been monitoring the hard-x-ray emission from the peculiar luminous star η Car since 1996 Feb. 9. Using a conversion factor of 3.22 × 10^{−12} erg s^{−1} cm^{−2} per PCA count, the 2–10-keV lightcurve (in PCU0-2) shows a significant long-term rise, with flux increasing from 6.8 × 10^{−11} to 1.0 × 10^{−10} erg s^{−1} cm^{−2} in the interval 1996 Dec.–1997 May. Highly significant, shorter-duration flaring has been observed on 1996 May 3, July 22, Oct. 24, 1997 Jan. 19, and Apr. 10, with peak fluxes of 6.5, 7.8, 7.3, 9.2, and 12.0 × 10^{−11} erg s^{−1} cm^{−2}, respectively, representing a flux increase of ∼ 9–40 percent over the non-flare emission. Timings of the flare peaks suggest a recurrence every 85.1 ± 5.4 days, with a duration of 20 days. The next flare event should occur in the interval 1997 June 23– July 13, with peak intensity predicted to occur on July 3. We encourage coordinated observations of η Car at all wavebands during this interval to determine if this is a persistent property and to search for the cause of this activity. Dates of upcoming and previous RXTE observations of η Car are available, along with the most recent RXTE x-ray lightcurve, from http://

lheawww.gsfc.nasa.gov/users/corcoran/eta car/eta car xte.html.”

SLX 1735–269 A. Bazzano, M. Cocchi, and P. Ubertini, Istituto di Astrofisica Spazi-

ale, CNR, Frascati; and J. Heise, J. in ’t Zand, and J. M. Muller, Space Research Organization of the Netherlands, Utrecht, report: “On Apr. 13.668 UT, an x-ray burst was detected with Wide Field Camera Number 1 on the BeppoSAX satellite from α = 17^{h}38^{m}17^{s}.8, δ = −27^{o}00′08′′ (equinox 2000.0, error circle radius 5′). The only known x-ray source within this error circle is SLX 1735–269. This is the first report of bursting behavior from this source, suggesting a neutron star as a likely source. The burst lasted 30 s and had a peak intensity of 900 mCrabs (2–8 keV). The steady emission of SLX 1735–269 had an average value of ∼ 10 mCrab during monitoring observations between Apr. 13.02 and 16.80.”

1997 May 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green